N7000 was shared by four people in ratio x:2:1:3. If the least was N702, what is x?
Found 3 solutions by Theo, greenestamps, ikleyn:
Answer by Theo(13342)

i get x = 3.971509972
as a fraction, that is equal to 1394/351.702
the ratio is therefore 1394/351:2:1:3
let k be the common factor.
you get:
1394/351*k + 2k + k + 3k = 7000
multiply both sides of this by 351 to get
1394k + 702k + 351k + 1053k = 2457000
combine like terms to get:
3500k = 2457000
solve for k to get:
k = 702
your ratio is therefore:
1394/351:2:1:3
x is therefore equal to 1394/351.
it's not an integer, but it is the right value for x.
multiply all of the factors by 351 to get:
1394:702:351:1053.
the common factor will now be (1394 + 702 + 351 + 1053) * k = 7000
solve for k to get:
k = 2
when x is an integer, the ratio becomes:
1394:702:351:1053 with k = 2.
k is the common factor.

your solution should be:

either x = 1394/351 when the ratio is 1394/351:2:1:3 and the common factor = k is 702, or x = 1394 when the ratio is 1394:702:351:1053 and the common factor = k is 2.

either ratio is good, depending on whether you want the ratio to be in whole number or you can allow fractions.

this is the best i can do with this.

let me know if you have any questions.

theo



Answer by greenestamps(13198)



In the given ratio, the "1" is the least share. So

the 2nd share is 2(N702)=N1404
the 3rd share is N702
the 4th share is 3(N702)=N2106

The first share is then N7000-(N1404+N702+N2106)=N2788

Then "x" in the problem is 2788/702=1394/351; which to several decimal places is 3.97151


Answer by ikleyn(52776)

.

It is really strange to me to see the number "702" in this problem as an input data.


            It is the same as if somebody would come stating that the duration
            of a day is  17  hours and  41  minutes, instead of  24  hours . . .


The number of 700, instead of 702, would be much more natural and adequate.

Twice, not, thrice check your source.